# Probabilistic Guarantees to Explicit Constructions: Local Properties of Linear Codes

### v1

@@ 5.3 Main Result
We now state and prove the central theorem of this section. Fix a reasonable $L$-LCL property $\mathcal{P}$, and $R_{\textrm{in}}\leq R_{\mathcal{P}}-\varepsilon$ for some $\varepsilon>0$. Let $\Delta=\varepsilon/2L$ be as defined above. Let $T_{\mathcal{P}}$ denote the maximum number of distinct matrices appearing in any local profile description $\mathcal{V}\in\mathcal{P}$. Let $G$ be the bipartite graph obtained from ˜3.4 by setting
@@
$\zeta=\frac{\delta_{\textrm{out}}}{2T_{\mathcal{P}}},$
@@
$d=O\left(\frac{1}{\zeta\eta^{2}}\right)=O\left(\frac{T_{\mathcal{P}}^{3}}{\Delta^{2}\cdot\delta_{\textrm{out}}}\right).$
For any integer $N$ satisfying $N>T_{\mathcal{P}}/\delta_{\textrm{out}}$, take $G$ to have $N$ vertices on both sides. Let $\mathcal{C}_{\textrm{in}}\subseteq{\mathbb{F}}_{q}^{d}$ be an ${\mathbb{F}}_{q}$-linear code of rate $R_{\textrm{in}}$, where $q$ is the minimum prime power guaranteed by ˜5.30, and let $\mathcal{C}_{\textrm{out}}\subseteq({\mathbb{F}}_{q}^{R_{\textrm{in}}d})^{N}$ be an ${\mathbb{F}}_{q}$-linear code having distance $\delta_{\textrm{out}}$. By definition, $\phi$ is now an ${\mathbb{F}}_{q}$-linear map of the form $\phi\colon{\mathbb{F}}_{q}^{R_{\textrm{in}}d}\rightarrow\mathcal{C}_{\textrm{in}}$.
@@
Let $\mathcal{C}_{\textrm{in}},\mathcal{C}_{\textrm{out}}$, and $G$ be as defined above. Furthermore, assume that $\mathcal{C}_{\textrm{in}}$ is a linear code whose existence is guaranteed by ˜5.30. Then, $\mathcal{C}_{\textrm{AEL}}(\mathcal{C}_{\textrm{in}},\mathcal{C}_{\textrm{out}},G)\subseteq{\mathbb{F}}_{Q}^{N}$ is an ${\mathbb{F}}_{q}$-linear code that does not satisfy $\mathcal{P}$.
@@
Since both $\mathcal{C}_{\textrm{out}}$ and the map $\phi$ are ${\mathbb{F}}_{q}$-linear, it follows that $\mathcal{C}_{\textrm{AEL}}$ is itself an ${\mathbb{F}}_{q}$-linear code.
Assume for contradiction that $\mathcal{C}_{\textrm{AEL}}$ satisfies $\mathcal{P}$. Then, there is a $\mathcal{V}=(f_{1},\mathbf{M}_{1}),\ldots,(f_{T},\mathbf{M}_{T}))\in\mathcal{P}$ such that $\mathcal{C}_{\textrm{AEL}}$ contains $\mathcal{V}$, which implies the existence of a matrix $A\in{\mathbb{F}}_{Q}^{N\times L}$, such that
@@
    $A\subseteq\mathcal{C}_{\textrm{AEL}}$,
@@
Let the columns of $A$ be pairwise distinct codewords $c_{1},\ldots,c_{L}\in\mathcal{C}_{\textrm{AEL}}\subseteq{\mathbb{F}}_{Q}^{N}$. Note that we can reinterpret these codewords as vectors in $({\mathbb{F}}_{q}^{d})^{N}$, and we choose to do so. We now perform three operations. Firstly, we create $A_{\mathsf{fl}}\in{\mathbb{F}}_{q}^{([N]\times[d])\times L}$ from $A$ by flattening every column of $A$ (see Definition˜3.3) and then collecting them in a matrix $A_{\mathsf{fl}}$. Therefore, for every $i\in[L]$, $A_{\mathsf{fl}}[][i]=(A[][i])_{\mathsf{fl}}$. Secondly, we create $A^{\mathsf{proj}}_{\mathsf{fl}}\in{\mathbb{F}}_{q}^{([N]\times[d])\times L}$ by performing the projection operation on $A_{\mathsf{fl}}$ (see Definition˜3.1).
We now instantiate the objects required to perform the third operation. As $R\leq R_{\mathcal{P}}-\varepsilon\leq R_{\mathcal{V}}-\varepsilon$, by ˜5.20 there exists a (canonical) subspace $W_{\mathcal{V}}$ for $\mathcal{V}$. In particular, because
@@
There exists a set of indices $L^{*}\subseteq[N]$ satisfying $\left|L^{*}\right|>(1-\delta_{\textrm{out}})N$ such that for every $\ell\in L^{*}$, there exists a $\mathcal{V}^{\mathsf{imp}}_{\Delta}\in\operatorname{Rob}_{\Delta}(\mathcal{V}^{\mathsf{imp}})$ such that the matrix $A^{\mathsf{proj}}_{\psi}(\ell)$ satisfies $\mathcal{V}^{\mathsf{imp}}_{\Delta}$.
@@
This immediately implies that one of the columns of $A^{\mathsf{proj}}_{\psi}$ is non-zero. Denote the index of this column by $e\in[L-\dim W_{\mathcal{V}}]$. Observe that the columns of $A^{\mathsf{proj}}_{\psi}$ are codewords in $\mathcal{C}_{\textrm{out}}\circ\mathcal{C}_{\textrm{in}}$. Indeed, the columns of $A_{\mathsf{fl}}$ are (flattened) codewords of $\mathcal{C}_{\textrm{AEL}}$, and so the columns of $A^{\mathsf{proj}}_{\mathsf{fl}}$ are codewords in $\mathcal{C}_{\textrm{out}}\circ\mathcal{C}_{\textrm{in}}$. It follows directly from the ${\mathbb{F}}_{q}$-linearity of $\mathcal{C}_{\textrm{out}}$ and $\mathcal{C}_{\textrm{in}}$ that $\mathcal{C}_{\textrm{out}}\circ\mathcal{C}_{\textrm{in}}$ is also ${\mathbb{F}}_{q}$-linear. Moreover, since $A^{\mathsf{proj}}_{\psi}$ is obtained by applying an ${\mathbb{F}}_{q}$-linear map to the rows of $A^{\mathsf{proj}}_{\mathsf{fl}}$, the columns of $A^{\mathsf{proj}}_{\psi}$ are ${\mathbb{F}}_{q}$-linear combinations of columns of $A^{\mathsf{proj}}_{\mathsf{fl}}$, and therefore are also codewords of $\mathcal{C}_{\textrm{out}}\circ\mathcal{C}_{\textrm{in}}$. In particular, this holds for the non-zero column $A^{\mathsf{proj}}_{\psi}[][e]$, and therefore is of the following form
@@
where $c$ is a codeword in $\mathcal{C}_{\textrm{out}}$. This in turn implies that at least $\delta_{\textrm{out}}N$ submatrices of the form $A^{\mathsf{proj}}_{\psi}(\ell)$ are non-zero, because the column indexed by $e$ is non-zero for all such submatrices. Denote this set of indices by $S\subseteq[N]$.
By ˜5.33, and the fact that $|S|\geq\delta_{\textrm{out}}N$, there is at least one index $\ell^{*}\in(S\cap L^{*})$. Since the columns of $A^{\mathsf{proj}}_{\psi}$ are codewords of $\mathcal{C}_{\textrm{out}}\circ\mathcal{C}_{\textrm{in}}$, it follows that every column of the submatrix $A^{\mathsf{proj}}_{\psi}(\ell^{*})$ is a codeword of $\mathcal{C}_{\textrm{in}}$. Therefore, $A^{\mathsf{proj}}_{\psi}(\ell^{*})$ has the following properties:
@@
    $A^{\mathsf{proj}}_{\psi}(\ell^{*})\subseteq\mathcal{C}_{\textrm{in}}$,
@@
This contradicts the fact that $\mathcal{C}_{\textrm{in}}$ is a code satisfying the guarantee mentioned in ˜5.30.
@@
$\left|\{\ell\in V_{L}\mid\left|\Gamma(\ell)\cap S_{t}\right|<(f_{t}-\eta)d\}\right|\leq\zeta N=\frac{\delta_{\textrm{out}}}{2T_{\mathcal{P}}}\cdot N<\frac{\delta_{\textrm{out}}}{T}\cdot N.$
@@
$\left|\{\ell\in V_{L}\mid\forall t\in T_{\beta},\left|\Gamma(\ell)\cap S_{t}\right|\geq\left(f_{t}-\eta\right)d\}\right|>(1-\delta_{\textrm{out}})N.$ (16)
Denote this set by $L^{*}\subseteq[N]$, and note that $\left|L^{*}\right|>(1-\delta_{\textrm{out}})N$.
@@
    $\leq\left(1-\sum_{t\in T_{\beta}}f_{t}\right)\cdot d+\frac{\Delta\cdot d}{4}+T_{\mathcal{P}}$
@@ 6 Consequences
For a reasonable $L$-LCL property $\mathcal{P}$, let $R_{\textrm{in}}=R_{\mathcal{P}}-\varepsilon$ for some $\varepsilon>0$. Let $\mathcal{C}_{\textrm{in}},\mathcal{C}_{\textrm{out}}$, and $G$ be as defined in Theorem˜5.32. Furthermore, let $\mathcal{C}_{\textrm{out}}\subseteq({\mathbb{F}}_{q}^{R_{\textrm{in}}d})^{N}$ be an ${\mathbb{F}}_{q}$-linear code with rate $R_{\textrm{out}}=1-\varepsilon$ and distance $\delta_{\textrm{out}}\geq\varepsilon^{3}$. Then, $\mathcal{C}_{\textrm{AEL}}(\mathcal{C}_{\textrm{in}},\mathcal{C}_{\textrm{out}},G)\subseteq{\mathbb{F}}_{Q}^{N}$ is an ${\mathbb{F}}_{q}$-linear code that does not satisfy $\mathcal{P}$, has rate $R_{\textrm{AEL}}>R_{\mathcal{P}}-2\varepsilon$, and
@@
Additionally, $\mathcal{C}_{\textrm{AEL}}(\mathcal{C}_{\textrm{in}},\mathcal{C}_{\textrm{out}},G)$ can be constructed in time ${\mathrm{poly}}(N)$.
@@
The ${\mathbb{F}}_{q}$-linearity of $\mathcal{C}_{\textrm{AEL}}$, and the fact that it does not satisfy $\mathcal{P}$ are established in Theorem˜5.32. The rate is given by
$R_{\textrm{AEL}}=\frac{\log_{Q}(\left|\mathcal{C}_{\textrm{AEL}}\right|)}{N}.$
Indeed, it is easy to see that $\left|\mathcal{C}_{\textrm{AEL}}\right|=\left|\mathcal{C}_{\textrm{out}}\right|=q^{R_{\textrm{in}}R_{\textrm{out}}Nd}$. Recalling that $Q=q^{d}$, a simple calculation gives $R_{\textrm{AEL}}=R_{\textrm{in}}\cdot R_{\textrm{out}}=(R_{\mathcal{P}}-\varepsilon)(1-\varepsilon)>R_{\mathcal{P}}-2\varepsilon$. The value for $d$ is obtained by recalling that $d=O(1/\zeta\eta^{2})$, $\eta=\Delta/(4T_{\mathcal{P}})$, $\zeta=\delta_{\textrm{out}}/2T_{\mathcal{P}},\Delta=\varepsilon/2L$, and plugging in the value for $\delta_{\textrm{out}}$ in $\zeta$.
We note that explicit constructions of ${\mathbb{F}}_{q}$-linear codes having rate $1-\varepsilon$ and distance $\varepsilon^{3}$ that are constructible in time ${\mathrm{poly}}(N)$ can be obtained by using Tanner codes (see Corollary 11.4.8 in [GRS23]). Moreover, the graph $G$ can be constructed in time ${\mathrm{poly}}(N)$, by ˜3.4. By Lemma˜5.31, $\mathcal{C}_{\textrm{in}}$ can be constructed in time $q^{{\mathrm{poly}}(d,L)}\leq{\mathrm{poly}}(N)$. Upon invoking ˜3.5, we see that $\mathcal{C}_{\textrm{AEL}}(\mathcal{C}_{\textrm{in}},\mathcal{C}_{\textrm{out}},G)$ can be constructed in time ${\mathrm{poly}}(N)$. ∎
We now present a general definition of list-recovery, from which the special cases of interest can be derived.

### v2

@@ 5.3 Main Result
We now state and prove the central theorem of this section. Fix an $L$-LCL property $\mathcal{P}$, and $R_{\textnormal{in}}\leq R_{\mathcal{P}}-\varepsilon$ for some $\varepsilon>0$. Let $\Delta=\varepsilon/2L$ be as defined above. Let $T_{\mathcal{P}}$ denote the maximum number of distinct matrices appearing in any local profile description $\mathcal{V}\in\mathcal{P}$. Let $G$ be the bipartite graph obtained from ˜3.4 by setting
@@
$\zeta=\frac{\delta_{\textnormal{out}}}{2T_{\mathcal{P}}},$
@@
$d=O\left(\frac{1}{\zeta\eta^{2}}\right)=O\left(\frac{T_{\mathcal{P}}^{3}}{\Delta^{2}\cdot\delta_{\textnormal{out}}}\right).$
For any integer $N$ satisfying $N>T_{\mathcal{P}}/\delta_{\textnormal{out}}$, take $G$ to have $N$ vertices on both sides. Let $\mathcal{C}_{\textnormal{in}}\subseteq{\mathbb{F}}_{q}^{d}$ be an ${\mathbb{F}}_{q}$-linear code of rate $R_{\textnormal{in}}$, where $q$ is the minimum prime power guaranteed by ˜5.30, and let $\mathcal{C}_{\textnormal{out}}\subseteq({\mathbb{F}}_{q}^{R_{\textnormal{in}}d})^{N}$ be an ${\mathbb{F}}_{q}$-linear code having distance $\delta_{\textnormal{out}}$. By definition, $\phi$ is now an ${\mathbb{F}}_{q}$-linear map of the form $\phi\colon{\mathbb{F}}_{q}^{R_{\textnormal{in}}d}\rightarrow\mathcal{C}_{\textnormal{in}}$.
@@
Let $\mathcal{C}_{\textnormal{in}},\mathcal{C}_{\textnormal{out}}$, and $G$ be as defined above. Furthermore, assume that $\mathcal{C}_{\textnormal{in}}$ is a linear code whose existence is guaranteed by ˜5.30. Then, $\mathcal{C}_{\textnormal{AEL}}(\mathcal{C}_{\textnormal{in}},\mathcal{C}_{\textnormal{out}},G)\subseteq{\mathbb{F}}_{Q}^{N}$ is an ${\mathbb{F}}_{q}$-linear code that does not satisfy $\mathcal{P}$.
@@
Since both $\mathcal{C}_{\textnormal{out}}$ and the map $\phi$ are ${\mathbb{F}}_{q}$-linear, it follows that $\mathcal{C}_{\textnormal{AEL}}$ is itself an ${\mathbb{F}}_{q}$-linear code.
Assume for contradiction that $\mathcal{C}_{\textnormal{AEL}}$ satisfies $\mathcal{P}$. Then, there is a $\mathcal{V}=(f_{1},\mathbf{M}_{1}),\ldots,(f_{T},\mathbf{M}_{T}))\in\mathcal{P}$ such that $\mathcal{C}_{\textnormal{AEL}}$ contains $\mathcal{V}$, which implies the existence of a matrix $A\in{\mathbb{F}}_{Q}^{N\times L}$, such that
@@
    $A\subseteq\mathcal{C}_{\textnormal{AEL}}$,
@@
Let the columns of $A$ be pairwise distinct codewords $c_{1},\ldots,c_{L}\in\mathcal{C}_{\textnormal{AEL}}\subseteq{\mathbb{F}}_{Q}^{N}$. Note that we can reinterpret these codewords as vectors in $({\mathbb{F}}_{q}^{d})^{N}$, and we choose to do so. We now perform three operations. Firstly, we create $A_{\mathsf{fl}}\in{\mathbb{F}}_{q}^{([N]\times[d])\times L}$ from $A$ by flattening every column of $A$ (see Definition˜3.3) and then collecting them in a matrix $A_{\mathsf{fl}}$. Therefore, for every $i\in[L]$, $A_{\mathsf{fl}}[][i]=(A[][i])_{\mathsf{fl}}$. Secondly, we create $A^{\mathsf{proj}}_{\mathsf{fl}}\in{\mathbb{F}}_{q}^{([N]\times[d])\times L}$ by performing the projection operation on $A_{\mathsf{fl}}$ (see Definition˜3.1).
We now instantiate the objects required to perform the third operation. As $R\leq R_{\mathcal{P}}-\varepsilon\leq R_{\mathcal{V}}-\varepsilon$, by ˜5.19 there exists a (canonical) subspace $W_{\mathcal{V}}$ for $\mathcal{V}$. In particular, because
@@
There exists a set of indices $L^{*}\subseteq[N]$ satisfying $\left|L^{*}\right|>(1-\delta_{\textnormal{out}})N$ such that for every $\ell\in L^{*}$, there exists a $\mathcal{V}^{\mathsf{imp}}_{\Delta}\in\operatorname{Rob}_{\Delta}(\mathcal{V}^{\mathsf{imp}})$ such that the matrix $A^{\mathsf{proj}}_{\psi}(\ell)$ satisfies $\mathcal{V}^{\mathsf{imp}}_{\Delta}$.
@@
This immediately implies that one of the columns of $A^{\mathsf{proj}}_{\psi}$ is non-zero. Denote the index of this column by $e\in[L-\dim W_{\mathcal{V}}]$. Observe that the columns of $A^{\mathsf{proj}}_{\psi}$ are codewords in $\mathcal{C}_{\textnormal{out}}\circ\mathcal{C}_{\textnormal{in}}$. Indeed, the columns of $A_{\mathsf{fl}}$ are (flattened) codewords of $\mathcal{C}_{\textnormal{AEL}}$, and so the columns of $A^{\mathsf{proj}}_{\mathsf{fl}}$ are codewords in $\mathcal{C}_{\textnormal{out}}\circ\mathcal{C}_{\textnormal{in}}$. It follows directly from the ${\mathbb{F}}_{q}$-linearity of $\mathcal{C}_{\textnormal{out}}$ and $\mathcal{C}_{\textnormal{in}}$ that $\mathcal{C}_{\textnormal{out}}\circ\mathcal{C}_{\textnormal{in}}$ is also ${\mathbb{F}}_{q}$-linear. Moreover, since $A^{\mathsf{proj}}_{\psi}$ is obtained by applying an ${\mathbb{F}}_{q}$-linear map to the rows of $A^{\mathsf{proj}}_{\mathsf{fl}}$, the columns of $A^{\mathsf{proj}}_{\psi}$ are ${\mathbb{F}}_{q}$-linear combinations of columns of $A^{\mathsf{proj}}_{\mathsf{fl}}$, and therefore are also codewords of $\mathcal{C}_{\textnormal{out}}\circ\mathcal{C}_{\textnormal{in}}$. In particular, this holds for the non-zero column $A^{\mathsf{proj}}_{\psi}[][e]$, and therefore is of the following form
@@
where $c$ is a codeword in $\mathcal{C}_{\textnormal{out}}$. This in turn implies that at least $\delta_{\textnormal{out}}N$ submatrices of the form $A^{\mathsf{proj}}_{\psi}(\ell)$ are non-zero, because the column indexed by $e$ is non-zero for all such submatrices. Denote this set of indices by $S\subseteq[N]$.
By ˜5.33, and the fact that $|S|\geq\delta_{\textnormal{out}}N$, there is at least one index $\ell^{*}\in(S\cap L^{*})$. Since the columns of $A^{\mathsf{proj}}_{\psi}$ are codewords of $\mathcal{C}_{\textnormal{out}}\circ\mathcal{C}_{\textnormal{in}}$, it follows that every column of the submatrix $A^{\mathsf{proj}}_{\psi}(\ell^{*})$ is a codeword of $\mathcal{C}_{\textnormal{in}}$. Therefore, $A^{\mathsf{proj}}_{\psi}(\ell^{*})$ has the following properties:
@@
    $A^{\mathsf{proj}}_{\psi}(\ell^{*})\subseteq\mathcal{C}_{\textnormal{in}}$,
@@
This contradicts the fact that $\mathcal{C}_{\textnormal{in}}$ is a code satisfying the guarantee mentioned in ˜5.30.
@@
$\left|\{\ell\in V_{L}\mid\left|\Gamma(\ell)\cap S_{t}\right|<(f_{t}-\eta)d\}\right|\leq\zeta N=\frac{\delta_{\textnormal{out}}}{2T_{\mathcal{P}}}\cdot N<\frac{\delta_{\textnormal{out}}}{T}\cdot N.$
@@
$\left|\{\ell\in V_{L}\mid\forall t\in T_{\beta},\left|\Gamma(\ell)\cap S_{t}\right|\geq\left(f_{t}-\eta\right)d\}\right|>(1-\delta_{\textnormal{out}})N.$ (16)
Denote this set by $L^{*}\subseteq[N]$, and note that $\left|L^{*}\right|>(1-\delta_{\textnormal{out}})N$.
@@
    $\leq\Bigl(1-\sum_{t\in T_{\beta}}f_{t}\Bigr)\cdot d+\frac{\Delta\cdot d}{4}+T_{\mathcal{P}}$
@@ 6 Consequences
For an $L$-LCL property $\mathcal{P}$, let $R_{\textnormal{in}}=R_{\mathcal{P}}-\varepsilon$ for some $\varepsilon>0$. Let $\mathcal{C}_{\textnormal{in}},\mathcal{C}_{\textnormal{out}}$, and $G$ be as defined in Theorem˜5.32. Furthermore, let $\mathcal{C}_{\textnormal{out}}\subseteq({\mathbb{F}}_{q}^{R_{\textnormal{in}}d})^{N}$ be an ${\mathbb{F}}_{q}$-linear code with rate $R_{\textnormal{out}}=1-\varepsilon$ and distance $\delta_{\textnormal{out}}\geq\varepsilon^{3}$. Then, $\mathcal{C}_{\textnormal{AEL}}(\mathcal{C}_{\textnormal{in}},\mathcal{C}_{\textnormal{out}},G)\subseteq{\mathbb{F}}_{Q}^{N}$ is an ${\mathbb{F}}_{q}$-linear LDPC code that does not satisfy $\mathcal{P}$, has rate $R_{\textnormal{AEL}}>R_{\mathcal{P}}-2\varepsilon$, and
@@
Additionally, $\mathcal{C}_{\textnormal{AEL}}(\mathcal{C}_{\textnormal{in}},\mathcal{C}_{\textnormal{out}},G)$ can be constructed in time ${\mathrm{poly}}(N)$.
@@
The ${\mathbb{F}}_{q}$-linearity of $\mathcal{C}_{\textnormal{AEL}}$, and the fact that it does not satisfy $\mathcal{P}$ are established in Theorem˜5.32. The rate is given by
$R_{\textnormal{AEL}}=\frac{\log_{Q}(\left|\mathcal{C}_{\textnormal{AEL}}\right|)}{N}.$
Indeed, it is easy to see that $\left|\mathcal{C}_{\textnormal{AEL}}\right|=\left|\mathcal{C}_{\textnormal{out}}\right|=q^{R_{\textnormal{in}}R_{\textnormal{out}}Nd}$. Recalling that $Q=q^{d}$, a simple calculation gives $R_{\textnormal{AEL}}=R_{\textnormal{in}}\cdot R_{\textnormal{out}}=(R_{\mathcal{P}}-\varepsilon)(1-\varepsilon)>R_{\mathcal{P}}-2\varepsilon$. The value for $d$ is obtained by recalling that $d=O(1/\zeta\eta^{2})$, $\eta=\Delta/(4T_{\mathcal{P}})$, $\zeta=\delta_{\textnormal{out}}/2T_{\mathcal{P}},\Delta=\varepsilon/2L$, and plugging in the value for $\delta_{\textnormal{out}}$ in $\zeta$.
We note that explicit constructions of ${\mathbb{F}}_{q}$-linear codes having rate $1-\varepsilon$ and distance $\varepsilon^{3}$ that are constructible in time ${\mathrm{poly}}(N)$ can be obtained by using Tanner codes (see Corollary 11.4.8 in [GRS23]). We note that Tanner codes are LDPC codes. Moreover, the graph $G$ can be constructed in time ${\mathrm{poly}}(N)$, by ˜3.4. By Lemma˜5.31, $\mathcal{C}_{\textnormal{in}}$ can be constructed in time $q^{{\mathrm{poly}}(d,L)}\leq{\mathrm{poly}}(N)$. Upon invoking ˜3.5, we see that $\mathcal{C}_{\textnormal{AEL}}(\mathcal{C}_{\textnormal{in}},\mathcal{C}_{\textnormal{out}},G)$ is an LDPC code that can be constructed in time ${\mathrm{poly}}(N)$. ∎
Our constructions also allow for combining multiple LCL properties. Given a list of $L$-LCL properties $\mathcal{P}_{1},\mathcal{P}_{2},\ldots$ where the constraints are defined over the same field $\mathbb{F}_{q_{0}}$, we can take the union of the local profile descriptions to form $\mathcal{P}$:
$\mathcal{P}:=\bigcup_{i}\mathcal{P}_{i}=\bigcup_{i}\left\{\mathcal{V}\mid\mathcal{V}\in\mathcal{P}_{i}\right\}.$
Since the set $\mathcal{P}$ is comprised of $L$-local profile descriptions, $\mathcal{P}$ is an $L$-LCL property as well. Thus, we can apply Corollary˜6.1 to $\mathcal{P}$ and obtain an explicit construction that satisfies the complement of $\mathcal{P}$. We record this discussion as the following corollary:
Corollary 6.2 (Explicit construction for combinations of LCL properties).
For $L$-LCL properties $\mathcal{P}_{1},\mathcal{P}_{2},\ldots$, define $\mathcal{P}$ as above. Then, the explicit construction detailed in Corollary˜6.1 for $\mathcal{P}$ avoids satisfying $\mathcal{P}_{1},\mathcal{P}_{2},\ldots$ simultaneously.
6.1 List Recovery
We present a general definition of list-recovery, from which the special cases of interest can be derived.